Find y' if xy^2 = yx^2

1
Expert's answer
2022-03-29T12:01:26-0400

"x^{y\u00b2}=y^{x\u00b2}"

Take natural log of both sides

"x\u00b2lny=y\u00b2lnx"


Differentiate both sides with respect to x

"2xlny+\frac{x²}{y}y'=(2ylnx)y'+\frac{y²}{x}"


Multiply through by "xy"

"2x\u00b2ylny+{x\u00b3}y'=(2y\u00b2xlnx)y'+y\u00b3"


"=> {x\u00b3}y'-(2y\u00b2xlnx)y'=y\u00b3-2x\u00b2ylny"


"=> (x\u00b3-2y\u00b2xlnx)y'=y\u00b3-2x\u00b2ylny"


"=> y'=\frac{y³-2x²ylny}{x³-2y²xlnx}"